mathslover (mathslover):

Let z be a non-real complex number lying on the circle |z| = 1. Then prove that : z = \(\frac{ 1 + i \tan ( \frac{ arg(z)}{2} )}{1 - i \tan ( \frac{arg(z)}{2})}\)

mathslover (mathslover):

wait

mathslover (mathslover):

\[\large{z = \frac{ 1 + i \tan ( \frac{ arg(z)}{2} )}{1 - i \tan ( \frac{arg(z)}{2})}}\]

OpenStudy (anonymous):

non-real complex? do you mean just imaginary?

Parth (parthkohli):

In which grade do you learn this?

mathslover (mathslover):

@ParthKohli in 11th grade

mathslover (mathslover):

I was able to prove by taking RHS and simplifying and simplifying it into such a form which was exactly equivalent to the LHS i.e. 'z' But , the book says this is not a method of proving infact it is a method of "verifying" ... :(

mathslover (mathslover):

:P Well lemme post what I think,,,, it will take time...

Parth (parthkohli):

(nothing to do here)

OpenStudy (jamesj):

First thing to do is rationalize the denominator. Multiply top and bottom by 1 + i . tan(Arg(z)/2)

OpenStudy (aravindg):

you can also do this by taking modulus of given equation

OpenStudy (aravindg):